OpenStudy (anonymous):

I need help setting up an equation i know the constraints variables and objective functions but i don't know how to set it up

OpenStudy (anonymous):

what function do you need?

OpenStudy (anonymous):

ok I'm trying to figure out this problem or at least how to write it out

OpenStudy (anonymous):

Here are the variables 128kbs Female 64Kbs for female for every female character it takes 5 fps for every 10 avatars for every male it takes 10 fps for every 10 avatars 12milliseconds for character look up figure out maximum number of avatars that can be displayed with out going over 10mb 10240kb keeping a loading time under 2 seconds 2000milleseconds while staying at 30 fps with no avatars 121 so you have 91 fps to use

OpenStudy (anonymous):

wow.. that is really hard

OpenStudy (anonymous):

I do not know where to start.....

OpenStudy (anonymous):

64kb for male 128x+64y= 12x=>2000 5x+10y= 10x+10Y=

OpenStudy (anonymous):

thats like my best guess of setting it up

OpenStudy (anonymous):

first the variables which are the avatars 128kb and 64kb then the constraints which are the time needed for them 5x+10y= 10x+10y= 12x+y=<=2000

OpenStudy (anonymous):

then the objective function which is the maximum number of avatars x+y-10240z-91f=?

OpenStudy (anonymous):

guess I'm stuck

OpenStudy (anonymous):

and some one told me finite math was easy

OpenStudy (anonymous):

hmmm I also do no know how to do this... I am reallysorry

OpenStudy (anonymous):

the math is easy but the setup is hard

OpenStudy (anonymous):

thank you for trying

OpenStudy (anonymous):

hmm.. listen first put constants on one side and variables on the other then it gets easy to solve an equation. i hope it helped you!\[3x + 4=10\] put the constants on one side and the variable on the other. hehehe thanks hope it helps u!